Physical Chemistry

Two inflated balloonsof volumes600ml and 1500ml respectively are taken .If the maximum volume of inner and outer balloons are 800ml and 1800ml respectively then which balloon will burst first on gradual heating

meghana p
6 Years agoGrade

1 Answer

Arun
6 Years ago
If we take for the inner balloon so, it has volume of 600 ml and temperature of 300k.
 
So, on applying Charles law we know that volume at constant pressure is directly proportional to its temperature. So, we will get 600/300 = 800/T2.
 
T2 = 800 x 300 / 600
 
So, the value of T2 = 400 K.
 
Now, for the outer balloon again applying Charles law we will get that volume is directly proportional to its temperature at constant pressure.
 
So, we will get that 1500/300 = 1800 / T2.
 
T2 = 1800 x 300 / 1500.
 
On solving we will get T2 = 360 K.
 
Since, the  outer balloon is having the break point temperature to be less so, the outer balloon will burst first.